OpenStudy (anonymous):

two spheres of the same density have the ratio of 4 to 9 in surface area. if the small sphere weighs 10kg what does the sphere weigh

OpenStudy (anonymous):

surface area = 4pi*r^2 \[\frac{4*\Pi*r_1}{4} = \frac{4*\Pi*r_2}{9}\]

OpenStudy (ujjwal):

the ratio of surface areas is 4:9 which ultimately gives you r1^2/r2^2 = 4/9 thus r1^3/r2^3 = 8/27 now use density=mass/volume and equalize densities of both balls finally, you will get m1/m2 = r1^3/r2^3 But we have r1^3/r2^3 = 8/27 so, your answer is 8/27

OpenStudy (ujjwal):

oh. sorry you have value of m1 put it in final result, you will get m2..

OpenStudy (anonymous):

Thank you !